# Pázmány Péter sétány 1/A, 1117 Budapest, Hungary

### v1

@@ 1 Introduction
In numerous cases it was possible to find a closed expression for the above mentioned overlaps valid for all Bethe states of the integrable spin chain, see e.g. Kristjansen:2024dnm , and this fact was attributed to the matrix product state being a discrete version of Zamolodchikov’s integrable boundary state Zamolodchikov:1989fp ; Piroli:2017sei . The overlap formulas for integrable matrix product states contain a universal factor which involves the Gaudin matrix of the Bethe eigenstate being considered. More precisely, the quantity which appears can be viewed as the super determinant of the Gaudin matrix Kristjansen:2020mhn . The appearance of this structure was observed already in overlap formulas involving simple two-site product states such as the Néel state in the XXZ spin chain Pozsgay_2014 ; Brockmann_2014odd ; Brockmann_2014 . In addition to the superdeterminant of the Gaudin matrix overlap formulas for integrable matrix product states contain a non-universal pre-factor which encodes the properties of the boundary state. For particular types of matrix product states the pre-factor has a completely factorized form where the factors take the form of Baxter $Q$-functions. These are states built from repeated two-site states, i.e. matrix product states which generalize the Néel state. For the XXZ chain and general for $GL(N)$ spin chains they were studied in respectively Gombor:2021uxz and Gombor:2021hmj ; Gombor:2023bez where closed overlap formulas were likewise found. Such states exist also for integrable super spin chains where the particular matrix product states correspond to field theoretical defects in the limit where the Nahm pole boundary conditions become standard Dirichlet boundary conditions Kristjansen:2020mhn ; Kristjansen:2020vbe ; Kristjansen:2021abc .
@@
This is a state of the integrable $SO(6)$ spin chain in the fundamental representation, hence $i_{1},\ldots i_{2J}\in\{1,2,\ldots,6\}$. The six matrices involved, when collected in a list $\bar{\omega}=\{\omega_{j}\}_{j=1}^{6}$ take the form
@@
There exists an easily applicable test to determine whether one can expect a closed overlap formula to exist for a given matrix product state, consisting in checking whether it is annihilated by all the parity odd conserved charges of its integrable spin chain host Piroli:2017sei . In practice, it is even sufficient to check if this criterion is fulfilled for the first in the series of odd charges Buhl-Mortensen:2015gfd ; deLeeuw:2016ofj . In order to actually derive the overlap formula a deeper understanding of the concept of boundary integrability is needed. In the same way as one needs the so-called RTT relation (to be made explicit in the next section) and not just the commutation relation for transfer matrices when one wants to explicitly determine the eigenstates and eigenvectors of an integrable spin chain, a relation which we will denote as the KT relation is needed to explicitly derive overlap formulas for the chain with a boundary. It turns out that the appropriate mathematical tool to deal with this type of relation is that of representation theory of twisted Yangians Pozsgay:2018dzs ; DeLeeuw:2019ohp . The information about the boundary is encoded in a $K$-matrix which has to fulfil the former KT relation, and a $K$-matrix with this property precisely constitutes a representation of a twisted Yangian Molev:1997wp . The relevance of twisted Yangians for integrable systems with boundaries was understood also in e.g. Doikou:2004hy ; MacKay:2004tc . Twisted Yangians have a co-module property which we will exploit to show that given one K-matrix which solves the KT relation, one can perform a dressing by the Yangian itself which gives rise to a novel solution of the KT relation. This dressing is a quantum analogue of the well-known dressing procedure of Zakharov:1973pp ; Zakharov:1980ty ; Harnad:1983we by means of which one can generate novel solutions of a classically integrable equation in terms of already known solutions and which has earlier been used in the AdS/CFT context for generating novel solutions of the classical string equations of motion Spradlin:2006wk ; Kalousios:2006xy ; Jevicki:2007pk . The overlap formula corresponding to the one-dimensional representation of the Yangian can be derived by existing methods (at lest for vanishing twist) as the corresponding boundary state constitutes a special example of a two-site product state for which the overlap formula was determined for any GL(N) spin chain in Gombor:2021hmj ; Gombor:2023bez . $K$-matrices which result from dressing the trivial representation give rise to overlaps which can directly be expressed in terms of eigenvalues of a single transfer matrix. In the present paper we will need to go beyond this type of $K$-matrices.
@@
Our paper is organized as follows. We begin in section 2 by introducing the tools of group theory and of integrability needed to study the integrable $\mathfrak{so}(6)$ spin chain with twists and inhomogeneities in its fundamental representation. Subsequently, in section 3 we introduce the $K$-matrix and the crucial KT relation which defines the concept of an integrable boundary state. Making use of the fact that a K-matrix which solves the KT relation constitutes a representation of the twisted Yangian $Y^{+}(4)$ we devise a quantum dressing procedure which allows us to generate a series of integrable matrix product states starting from the trivial representation. In section 4.1 we turn to deriving, by means of examples, the generalized dressing formulas necessary to reach the two classes of integrable matrix product states which correspond to evaluation representations of $Y^{+}(4)$, the representations of relevance for the D3-D5 domain wall set-up. The details of of this derivation is relegated to appendices B and C whereas appendix A contains the explicit $K$-matrices of the evaluation representations. Finally, in section 5 by considering the limit of vanishing twist and impurities we recover the overlap formula for the general D3-D5 matrix product state given in eqn. (1.1), obtained numerically in DeLeeuw:2018cal . We also show that our dressing formulas are in fact valid for any representation of the quantum space. Section 6 contains our conclusion.
@@ 2.1 The $\mathfrak{gl}_{4}$ Lie-algebra and its representations
Let $e_{i}$ be the usual basis in $\mathbb{C}^{4}$ and let $e_{i,j}$ be the unit matrices in $\mathbb{C}^{4}$ defined by $e_{i,j}e_{k}=\delta_{j,k}e_{i}$. Furthermore, let $e_{[i,j]}$ for $1\leq i<j\leq 4$ be a basis of $\mathbb{C}^{6}$ coming from the anti-symmetrization of $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$ and let $e_{[i,j],[k,l]}$ be the unit matrices in $\mathbb{C}^{6}$ as $e_{[i,j],[k,l]}e_{[a,b]}=\delta_{k,a}\delta_{l,b}e_{[i,j]}$. We can identify the six dimensional basis with the three complex fields in ${\cal N}=4$ SYM and their complex conjugates as follows
@@
The matrices $e_{i,j}$ constitute the defining representation. Let us introduce the matrices $\mathcal{E}_{i,j}$ as
@@
which generate the 6 dimensional representation. For every 4-tuple $\lambda\equiv(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ we can define a highest weight irreducible representation (irrep) $E_{i,j}^{\lambda}$ of $\mathfrak{gl}_{4}$ as
@@
where $|0_{\lambda}\rangle$ is the highest weight state. We will make use of, he Gelfand-Tsetlin basis of the $\mathfrak{gl}_{4}$ irreps, see e.g. molev2002gelfandtsetlinbasesclassicallie . Every basis vector $|\Lambda\rangle$ of the irrep $E_{i,j}^{\lambda}$ corresponds to a GT-pattern
@@ 2.2 Transfer matrices
The first factor in the product space is referred to as the auxiliary space and the second one as the quantum space. The transfer matrices above satisfy the following algebra
@@
Figure 1: A graphical representation of the monodromy matrices and the RTT-relations. The red and blue lines correspond to the auxiliary space ($i,j=1,\dots 4$). The black lines are the 6 dimensional representations. The intersection of two red lines represents the R-matrix, $R(u)$, and the intersection of a red and a blue line the crossed R-matrix, $\bar{R}(u)$.
@@
$\tau_{\lambda}(u)$ $=\sum_{\Lambda}\left[\prod_{j=1}^{4}z_{j}^{\omega_{j}}\right]\frac{Q_{\theta}(u+1/2+\lambda_{2,1})Q_{\theta}(u-1/2+\lambda_{2,2})}{Q_{\theta}(u-1/2)}\prod_{k=1}^{3}\mathcal{F}_{\Lambda}^{(k)}(u),$ (2.23)
$\mathcal{F}_{\Lambda}^{(k)}(u)$ $=\frac{Q_{k}(u+\lambda_{k+1,k+1}-k/2)}{Q_{k}(u+\lambda_{k,k}-k/2)}\prod_{j=1}^{k}\frac{Q_{k}(u+\lambda_{k+1,j}+k/2-j+1)}{Q_{k}(u+\lambda_{k,j}+k/2-j+1)}\prod_{j=1}^{k-1}\frac{Q_{k}(u+\lambda_{k-1,j}+k/2-j)}{Q_{k}(u+\lambda_{k,j}+k/2-j)}.$
@@ 3 Integrable Boundary states and Twisted Yangians
Our goal is to calculate overlaps, $\langle\mbox{MPS}|\bar{u}\rangle$, between integrable matrix product states and on-shell Bethe eigenstates where the on-shell Bethe states were defined in the previous section. The purpose of this section is to define the integrable MPS’s. We first define integrable boundary states and corresponding K-matrices using the KT-relation Gombor:2021hmj ; Gombor:2023bez . The trace of the boundary state in the boundary space then gives the MPS. $K$-matrices which fulfil the KT relation constitute representations of the twisted Yangian $Y^{+}(4)$. We briefly review the properties of the evaluation representations of $Y^{+}(4)$ Molev:1997wp and introduce the corresponding MPS’s. These are the types of MPS’s which appear in the D3-D5 domain wall problem.
@@ 3.1 Definition of the KT-relation
where $i,j,k=1,\dots,4$ and refer to the auxiliary space. The indices $\alpha,\beta,\gamma$ take values in the set $\{1,\ldots,d_{B}\}$, and refer to a boundary vector space $\mathcal{V}_{B}\cong\mathbb{C}^{d_{B}}$ of dimension $d_{B}$, denoted as the bond dimension. The states $\langle\Psi_{\alpha,\beta}|$ are covectors in the quantum space, $\mathcal{H}$, and $K_{i,k}^{\alpha,\gamma}(u)$’s are scalars in this space.
We can thus define matrices in the boundary space (i.e. elements in $\mathrm{End}(\mathcal{V}_{B})$) associated to the two latter objects in the following way $\mathbf{K}_{i,j}(u)=\left\{K_{i,k}^{\alpha,\beta}(u)\right\}_{\alpha,\beta=1}^{d_{B}}\in\mathrm{End}(\mathcal{V}_{B})$ and $\langle\Psi|=\left\{\Psi_{\alpha,\beta}\right\}_{\alpha,\beta=1}^{d_{B}}\in\mathcal{H}^{*}\otimes\mathrm{End}(\mathcal{V}_{B})$. We can also introduce the notation $\mathbf{K}(u)=\sum_{i,j}e_{i,j}\otimes\mathbf{K}_{i,j}(u)$ for which the KT-relation can be written in a more compact form
@@
Figure 2: The graphical presentation of the K-matrices, boundary states. The red and blue lines correspond to the auxiliary space ($i,j=1,\dots 4$). The black lines are the 6-dimensional representations ($a,b=1,\dots 6$). The boundary space is denoted by orange lines ($\alpha,\beta=1,\dots d_{B}$).
Figure 3: The graphical presentation of the KT-relations. The red and blue lines correspond to the auxiliary space. The black lines are the 6 dimensional representations. The boundary space denoted by orange lines.
@@ 3.3 Twisted Yangians
The crossed reflection equation with this asymptotic expansion defines the twisted Yangian $Y^{+}(4)$ algebra. From the reflection equation we could derive that the operators $F_{i,j}$ generates an $\mathfrak{o}_{4}$ algebra
@@
The solutions of this equation can be obtained from the series expansion of the reflection equation Gombor:2019bun . The solution is
@@
for $i,j=1,\dots,4$. Therefore the symmetry algebra is $\mathfrak{o}_{4}$. From the condition (3.5) we can see that the twist is compatible with the boundary state only if the twist matrix is orthogonal, i.e., $G\in SO(4)$. The most general diagonal twist is

### v2

@@ 1 Introduction
In numerous cases it was possible to find a closed expression for the above mentioned overlaps valid for all Bethe states of the integrable spin chain, see e.g. Kristjansen:2024dnm , and this fact was attributed to the matrix product state being a discrete version of Zamolodchikov’s integrable boundary state Zamolodchikov:1989fp ; Piroli:2017sei . The overlap formulas for integrable matrix product states contain a universal factor which involves the Gaudin matrix of the Bethe eigenstate being considered. More precisely, the quantity which appears can be viewed as the super determinant of the Gaudin matrix Kristjansen:2020mhn . The appearance of this structure was observed already in overlap formulas involving simple two-site product states such as the Néel state in the XXZ spin chain Pozsgay_2014 ; Brockmann_2014odd ; Brockmann_2014 . In addition to the superdeterminant of the Gaudin matrix overlap formulas for integrable matrix product states contain a non-universal pre-factor which encodes the properties of the boundary state. For particular types of matrix product states the pre-factor has a completely factorized form where the factors take the form of Baxter $Q$-functions. These are states built from repeated two-site states, i.e. matrix product states which generalize the Néel state. For the XXZ chain and for general $GL(N)$ spin chains they were studied in respectively Gombor:2021uxz and Gombor:2021hmj ; Gombor:2023bez where closed overlap formulas were likewise found. Such states exist also for integrable super spin chains where the particular matrix product states correspond to field theoretical defects in the limit where the Nahm pole boundary conditions become standard Dirichlet boundary conditions Kristjansen:2020mhn ; Kristjansen:2020vbe ; Kristjansen:2021abc .
@@
This is a state of the integrable $SO(6)$ spin chain in the fundamental representation, hence $i_{1},\ldots,i_{2J}\in\{1,2,\ldots,6\}$. The six matrices involved, when collected in a list $\bar{\omega}=\{\omega_{j}\}_{j=1}^{6}$ take the form
@@
There exists an easily applicable test to determine whether one can expect a closed overlap formula to exist for a given matrix product state, consisting in checking whether it is annihilated by all the parity odd conserved charges of its integrable spin chain host Piroli:2017sei . In practice, it is even sufficient to check if this criterion is fulfilled for the first in the series of odd charges Buhl-Mortensen:2015gfd ; deLeeuw:2016ofj . In order to actually derive the overlap formula a deeper understanding of the concept of boundary integrability is needed. In the same way as one needs the so-called RTT relation (to be made explicit in the next section) and not just the commutation relation for transfer matrices when one wants to explicitly determine the eigenstates and eigenvectors of an integrable spin chain, a relation which we will denote as the KT relation is needed to explicitly derive overlap formulas for the chain with a boundary. It turns out that the appropriate mathematical tool to deal with this type of relation is that of representation theory of twisted Yangians Pozsgay:2018dzs ; DeLeeuw:2019ohp . The information about the boundary is encoded in a $K$-matrix which has to fulfil the former KT relation, and a $K$-matrix with this property precisely constitutes a representation of a twisted Yangian Molev:1997wp . The relevance of twisted Yangians for integrable systems with boundaries was understood also in e.g. Doikou:2004hy ; MacKay:2004tc . Twisted Yangians have a co-module property which we will exploit to show that given one $K$-matrix which solves the KT relation, one can perform a dressing by the Yangian itself which gives rise to a novel solution of the KT relation. This dressing is a quantum analogue of the well-known dressing procedure of Zakharov:1973pp ; Zakharov:1980ty ; Harnad:1983we by means of which one can generate novel solutions of a classically integrable equation in terms of already known solutions and which has earlier been used in the AdS/CFT context for generating novel solutions of the classical string equations of motion Spradlin:2006wk ; Kalousios:2006xy ; Jevicki:2007pk . The overlap formula corresponding to the one-dimensional representation of the Yangian can be derived by existing methods (at least for vanishing twist) as the corresponding boundary state constitutes a special example of a two-site product state for which the overlap formula was determined for any GL(N) spin chain in Gombor:2021hmj ; Gombor:2023bez . $K$-matrices which result from dressing the trivial representation give rise to overlaps which can directly be expressed in terms of eigenvalues of a single transfer matrix. In the present paper we will need to go beyond this type of $K$-matrices.
@@
Our paper is organized as follows. We begin in section 2 by introducing the tools of group theory and of integrability needed to study the integrable $\mathfrak{so}(6)$ spin chain with twists and inhomogeneities in its fundamental representation. Subsequently, in section 3 we introduce the $K$-matrix and the crucial KT relation which defines the concept of an integrable boundary state. Making use of the fact that a $K$-matrix which solves the KT relation constitutes a representation of the twisted Yangian $Y^{+}(4)$ we devise a quantum dressing procedure which allows us to generate a series of integrable matrix product states starting from the trivial representation. In section 4 we turn to deriving, by means of examples, the generalized dressing formulas necessary to reach the two classes of integrable matrix product states which correspond to evaluation representations of $Y^{+}(4)$, the representations of relevance for the D3-D5 domain wall set-up. The details of of this derivation are relegated to appendices B and C whereas appendix A contains the explicit $K$-matrices of the evaluation representations. Finally, in section 5 by considering the limit of vanishing twist and impurities we recover the overlap formula for the general D3-D5 matrix product state given in eqn. (1.1), obtained numerically in DeLeeuw:2018cal . We also show that our dressing formulas are in fact valid for any representation of the quantum space. Section 6 contains our conclusion.
@@ 2.1 The $\mathfrak{gl}_{4}$ Lie-algebra and its representations
Let $e_{i}$ be the usual basis in $\mathbb{C}^{4}$ and let $e_{i,j}$ be the unit matrices in $\mathbb{C}^{4}$ defined by $e_{i,j}e_{k}=\delta_{j,k}e_{i}$. Furthermore, let $e_{[i,j]}$ for $1\leq i<j\leq 4$ be a basis of $\mathbb{C}^{6}$ coming from the anti-symmetrization of $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$ and let $e_{[i,j],[k,l]}$ be the unit matrices in $\mathbb{C}^{6}$ which fulfil $e_{[i,j],[k,l]}e_{[a,b]}=\delta_{k,a}\delta_{l,b}e_{[i,j]}$. We can identify the six dimensional basis with the three complex fields in ${\cal N}=4$ SYM and their complex conjugates as follows
@@
The matrices $e_{i,j}$ constitute the defining representation. Let us introduce the matrices $\mathcal{E}_{i,j}$
@@
which generate the six-dimensional representation. For every 4-tuple $\lambda\equiv(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ we can define a highest weight irreducible representation (irrep) $E_{i,j}^{\lambda}$ of $\mathfrak{gl}_{4}$ by
@@
where $|0_{\lambda}\rangle$ is the highest weight state. We will make use of the Gelfand-Tsetlin basis of the $\mathfrak{gl}_{4}$ irreps, see e.g. molev2002gelfandtsetlinbasesclassicallie . Every basis vector $|\Lambda\rangle$ of the irrep $E_{i,j}^{\lambda}$ corresponds to a GT-pattern
@@ 2.2 Transfer matrices
The first factor in the product space is referred to as the auxiliary space and the second one as the quantum space. The Lax matrices above satisfy the following algebra
@@
Figure 1: A graphical representation of the monodromy matrices and the RTT-relations. The red and blue lines correspond to the auxiliary space ($i,j=1,\dots 4$). The black lines are the six-dimensional representations. The intersection of two red lines represents the R-matrix, $R(u)$, and the intersection of a red and a blue line the crossed R-matrix, $\bar{R}(u)$.
@@
$\tau_{\lambda}(u|\bar{u})$ $=\sum_{\Lambda}\left[\prod_{j=1}^{4}z_{j}^{\omega_{j}}\right]\frac{Q_{\theta}(u+1/2+\lambda_{2,1})Q_{\theta}(u-1/2+\lambda_{2,2})}{Q_{\theta}(u-1/2)}\prod_{k=1}^{3}\mathcal{F}_{\Lambda}^{(k)}(u|\bar{u}),$ (2.23)
$\mathcal{F}_{\Lambda}^{(k)}(u|\bar{u})$ $=\frac{Q_{k}(u+\lambda_{k+1,k+1}-k/2)}{Q_{k}(u+\lambda_{k,k}-k/2)}\prod_{j=1}^{k}\frac{Q_{k}(u+\lambda_{k+1,j}+k/2-j+1)}{Q_{k}(u+\lambda_{k,j}+k/2-j+1)}\prod_{j=1}^{k-1}\frac{Q_{k}(u+\lambda_{k-1,j}+k/2-j)}{Q_{k}(u+\lambda_{k,j}+k/2-j)}.$
@@ 3 Integrable Boundary states and Twisted Yangians
Our goal is to calculate overlaps, $\langle\mbox{MPS}|\bar{u}\rangle$, between integrable matrix product states and on-shell Bethe eigenstates where the on-shell Bethe states were defined in the previous section. The purpose of this section is to define the integrable MPSs. We first define integrable boundary states and corresponding $K$-matrices using the KT-relation Gombor:2021hmj ; Gombor:2023bez . The trace of the boundary state in the boundary space then gives the MPS. $K$-matrices which fulfil the KT relation constitute representations of the twisted Yangian $Y^{+}(4)$. We briefly review the properties of the evaluation representations of $Y^{+}(4)$ Molev:1997wp and introduce the corresponding MPSs. These are the types of MPSs which appear in the D3-D5 domain wall problem.
@@ 3.1 Definition of the KT-relation
where $i,j,k=1,\dots,4$ and refer to the auxiliary space. The indices $\alpha,\beta,\gamma$ take values in the set $\{1,\ldots,d_{B}\}$, and refer to a boundary vector space $\mathcal{V}_{B}\cong\mathbb{C}^{d_{B}}$ of dimension $d_{B}$, denoted as the bond dimension. The states $\langle\Psi_{\alpha,\beta}|$ are covectors in the quantum space, $\mathcal{H}$, and $K_{i,k}^{\alpha,\gamma}(u)$’s are scalars in this space. We can thus define matrices in the boundary space (i.e. elements in $\mathrm{End}(\mathcal{V}_{B})$) associated to the two latter objects in the following way $\mathbf{K}_{i,j}(u)=\left\{K_{i,k}^{\alpha,\beta}(u)\right\}_{\alpha,\beta=1}^{d_{B}}\in\mathrm{End}(\mathcal{V}_{B})$ and $\langle\Psi|=\left\{\Psi_{\alpha,\beta}\right\}_{\alpha,\beta=1}^{d_{B}}\in\mathcal{H}^{*}\otimes\mathrm{End}(\mathcal{V}_{B})$. We can also introduce the notation $\mathbf{K}(u)=\sum_{i,j}e_{i,j}\otimes\mathbf{K}_{i,j}(u)$ for which the KT-relation can be written in a more compact form
@@
Figure 2: The graphical presentation of the $K$-matrix and the boundary state. The red and blue lines correspond to the auxiliary space ($i,j=1,\dots 4$). The black lines are the six-dimensional representations ($a,b=1,\dots 6$). The boundary space is denoted by orange lines ($\alpha,\beta=1,\dots,d_{B}$).
Figure 3: The graphical presentation of the KT-relation. The red and blue lines correspond to the auxiliary space. The black lines are the six-dimensional representations. The boundary space is denoted by orange lines.
@@ 3.3 Twisted Yangians
The crossed reflection equation with this asymptotic expansion defines the twisted Yangian $Y^{+}(4)$ algebra. From the reflection equation we could derive that the operators $F_{i,j}$ generate an $\mathfrak{o}_{4}$ algebra
@@
The solution of this equation can be obtained from the series expansion of the reflection equation Gombor:2019bun which leads to
@@
for $i,j=1,\dots,4$. Therefore the symmetry algebra is $\mathfrak{o}_{4}$. From the condition (3.5) we see that the twist is compatible with the boundary state only if the twist matrix is orthogonal, i.e. $G\in SO(4)$. The most general diagonal twist is
@@
3.4 Selection rules for on-shell overlaps
We can apply the integrability condition (3.6) to an on-shell Bethe state and get
$(\tau(v|\bar{u})-\hat{\tau}(-v|\bar{u}))\langle\mathrm{MPS}|\bar{u}\rangle=0$ (3.18)
Non-vanishing on-shell overlaps $\langle\mathrm{MPS}|\bar{u}\rangle\neq 0$ require the condition
$\tau(v|\bar{u})=\hat{\tau}(-v|\bar{u}).$ (3.19)
The transfer matrix eigenvalues can be read off from eqn. (2.23) and take the form
$\begin{split}\tau(v|\bar{u})&=z_{1}\frac{Q_{\theta}(v+3/2)}{Q_{\theta}(v+1/2)}\frac{Q_{1}(v-1/2)}{Q_{1}(v+1/2)}+z_{2}\frac{Q_{\theta}(v+3/2)}{Q_{\theta}(v+1/2)}\frac{Q_{1}(v+3/2)}{Q_{1}(v+1/2)}\frac{Q_{2}(v)}{Q_{2}(v+1)}+\\ &+z_{3}\frac{Q_{2}(v+2)}{Q_{2}(v+1)}\frac{Q_{3}(v+1/2)}{Q_{3}(v+3/2)}+z_{4}\frac{Q_{3}(v+5/2)}{Q_{3}(v+3/2)},\\ \hat{\tau}(v|\bar{u})&=z_{4}^{-1}\frac{Q_{3}(v-5/2)}{Q_{3}(v-3/2)}+z_{3}^{-1}\frac{Q_{2}(v-2)}{Q_{2}(v-1)}\frac{Q_{3}(v-1/2)}{Q_{3}(v-3/2)}+\\ &+z_{2}^{-1}\frac{Q_{\theta}(v-3/2)}{Q_{\theta}(v-1/2)}\frac{Q_{1}(v-3/2)}{Q_{1}(v-1/2)}\frac{Q_{2}(v)}{Q_{2}(v-1)}+z_{1}^{-1}\frac{Q_{\theta}(v-3/2)}{Q_{\theta}(v-1/2)}\frac{Q_{1}(v+1/2)}{Q_{1}(v-1/2)}.\end{split}$ (3.20)
For the untwisted case ($z_{j}=1$) we can see that the condition (3.19) is satisfied when
$Q_{k}(-v)=(-1)^{n_{k}}Q_{k}(v),$ (3.21)
for $k=1,2,3$ which is the well-known pairing condition for the roots DeLeeuw:2018cal ; Gombor:2020kgu .
In the twisted case, however, it is not possible to formulate the selection rule in a nice way with these Bethe roots and $Q$-functions. To obtain the selection rules in a compact form we need to introduce the full system of $\mathcal{Q}$-functions Tsuboi:1998ne ( $\mathcal{Q}_{j},\mathcal{Q}_{jk},\mathcal{Q}_{jkl}$ where $j,k,l\in\{1,2,3,4\}$), for reviews see Gromov:2019icz ; Kazakov:2018ugh ; Levkovich-Maslyuk:2019awk . The relevant $\mathcal{Q}\mathcal{Q}$-relations read
$\begin{split}\mathcal{Q}_{j}(v+1/2)\mathcal{Q}_{k}(v-1/2)-\mathcal{Q}_{j}(v-1/2)\mathcal{Q}_{k}(v+1/2)&\sim\mathcal{Q}_{jk}(v),\\ \mathcal{Q}_{jk}(v+1/2)\mathcal{Q}_{jl}(v-1/2)-\mathcal{Q}_{jk}(v-1/2)\mathcal{Q}_{jl}(v+1/2)&\sim Q_{\theta}(v)\mathcal{Q}_{j}(v)\mathcal{Q}_{jkl}(v),\\ \mathcal{Q}_{jkl}(v+1/2)\mathcal{Q}_{jkn}(v-1/2)-\mathcal{Q}_{jkl}(v-1/2)\mathcal{Q}_{jkn}(v+1/2)&\sim\mathcal{Q}_{jk}(v).\end{split}$ (3.22)
We can indentify the previous $Q$-functions in terms of the $\mathcal{Q}$-functions as $\mathcal{Q}_{1}(u)=z_{1}^{-u}Q_{1}(u)$, $\mathcal{Q}_{12}(u)=z_{1}^{-u}z_{2}^{-u}Q_{2}(u)$ and $\mathcal{Q}_{123}(u)=z_{1}^{-u}z_{2}^{-u}z_{3}^{-u}Q_{3}(u)$. The transfer matrix eigenvalues can be expressed as
$\begin{split}\tau(v|\bar{u})&=\frac{Q_{\theta}(v+3/2)}{Q_{\theta}(v+1/2)}\frac{\mathcal{Q}_{j}(v-1/2)}{\mathcal{Q}_{j}(v+1/2)}+\frac{Q_{\theta}(v+3/2)}{Q_{\theta}(v+1/2)}\frac{\mathcal{Q}_{j}(v+3/2)}{\mathcal{Q}_{j}(v+1/2)}\frac{\mathcal{Q}_{jk}(v)}{\mathcal{Q}_{jk}(v+1)}+\\ &+\frac{\mathcal{Q}_{jk}(v+2)}{\mathcal{Q}_{jk}(v+1)}\frac{\mathcal{Q}_{jkl}(v+1/2)}{\mathcal{Q}_{jkl}(v+3/2)}+\frac{\mathcal{Q}_{jkl}(v+5/2)}{\mathcal{Q}_{jkl}(v+3/2)},\\ \hat{\tau}(v|\bar{u})&=\frac{\mathcal{Q}_{jkl}(v-5/2)}{\mathcal{Q}_{jkl}(v-3/2)}+\frac{\mathcal{Q}_{jk}(v-2)}{\mathcal{Q}_{jk}(v-1)}\frac{\mathcal{Q}_{jkl}(v-1/2)}{\mathcal{Q}_{jkl}(v-3/2)}+\\ &+\frac{Q_{\theta}(v-3/2)}{Q_{\theta}(v-1/2)}\frac{\mathcal{Q}_{j}(v-3/2)}{\mathcal{Q}_{j}(v-1/2)}\frac{\mathcal{Q}_{jk}(v)}{\mathcal{Q}_{jk}(v-1)}+\frac{Q_{\theta}(v-3/2)}{Q_{\theta}(v-1/2)}\frac{\mathcal{Q}_{j}(v+1/2)}{\mathcal{Q}_{j}(v-1/2)},\end{split}$ (3.23)
where we assumed that $z_{1}z_{2}z_{3}z_{4}=1$. We already saw that the boundary state is compatible with the twist when $z_{j}^{-1}=z_{5-j}$. We see that the condition (3.19) leads to the selection rule
$\mathcal{Q}_{k}(-v)=\pm\mathcal{Q}_{5-k}(v),$ (3.24)
for $k=1,2,3,4$.